# Hier-SLAM++: Neuro-Symbolic Semantic SLAM with a Hierarchically Categorical Gaussian Splatting

### v1

@@ Abstract
We propose Hier-SLAM++, a comprehensive Neuro-Symbolic semantic 3D Gaussian Splatting SLAM method with both RGB-D and monocular input featuring an advanced hierarchical categorical representation, which enables accurate pose estimation as well as global 3D semantic mapping. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making scene understanding particularly challenging and costly. To address this problem, we introduce a novel and general hierarchical representation that encodes both semantic and geometric information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs) as well as the 3D generative model. By utilizing the proposed hierarchical tree structure, semantic information is symbolically represented and learned in an end-to-end manner. We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Additionally, we propose an improved SLAM system to support both RGB-D and monocular inputs using a feed-forward model. To the best of our knowledge, this is the first semantic monocular Gaussian Splatting SLAM system, significantly reducing sensor requirements for 3D semantic understanding and broadening the applicability of semantic Gaussian SLAM system. We conduct experiments on both synthetic and real-world datasets, demonstrating superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM systems, while significantly reducing storage and training time requirements.
@@ I Introduction
Visual Simultaneous Localization and Mapping (SLAM) is a key technology for ego-motion estimation and scene perception, widely employed in applications across various domains, such as autonomous drones [1], self-driving vehicles [2], and interactive experiences in Augmented Reality (AR) and Virtual Reality (VR) [3]. Semantic information, which provides high-level knowledge about an environment, is indispensable for comprehensive scene understanding. Thus, it plays a crucial role in enabling intelligent robots to perform various complex tasks. In recent decades, advances in image segmentation and enhanced map representations have accelerated progress in semantic visual SLAM [4, 5, 6, 7, 8, 9, 10, 11, 7, 12].
@@
Recently, 3D Gaussian Splatting has gained attention as a leading 3D world representation method [13, 14, 15], thanks to its fast rendering and optimization efficiency enabled by the highly parallelized rasterization of 3D primitives. Specifically, 3D Gaussian Splatting models the continuous distributions of geometric parameters through Gaussian distributions. This approach not only improves the quality of 3D representation but also supports efficient optimization, making it particularly well-suited for SLAM tasks. While promising, current SLAM systems based on 3D Gaussian Splatting [16, 17, 18, 19, 20] primarily focus on geometric reconstruction. However, the lack of semantic integration limits its potential for complex tasks such as visual navigation, planning, and autonomous driving. Addressing this gap by incorporating semantic understanding without sacrificing efficiency remains a key challenge.
A naive solution would be to augment 3D primitives with discrete semantic labels and parameterize their distribution using a categorical distribution, i.e., a flat Softmax-based representation. However, 3D Gaussian Splatting is already a storage-intensive approach [21, 22], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would further increase storage requirements and processing time, with grow linear in the number of semantic classes, making it particularly impractical for complex scene understanding. Furthermore, given a fixed number of observations (video frames), increasing the number of optimizable parameters leads to the curse of dimensionality, resulting in degraded optimization performance in high-dimensional spaces, such as when handling a large number of semantic classes. Overall, this added overhead for semantic learning is especially problematic for robotics applications, where real-time performance and limited resources are critical. Recent studies have explored alternative methods to mitigate these challenges. For instance, [23] directly learns a 3-channel RGB visualization to avoid semantic labels learning. [24] adopts a flat semantic representation and relies on a pre-trained foundation model to generate an embedding feature map rather than an explicit 3D semantic label map.
Unlike flat representations, semantic information naturally organizes into a hierarchical structure of classes, as shown in Fig. 1. The attributes (such as structures) and properties (such as surfaces) of the semantic classes are extracted as symbolic nodes to generate a hierarchical tree, where each semantic class is represented as a hierarchical expression based on a root-to-leaf path. This hierarchy enable efficient coding of extensive information with a relatively small number of symbolic nodes to learn, resulting in a compact representation. For example, a binary tree with a depth of $10$ can represent $2^{10}$ classes, allowing for the encoding of 1,024 classes using only 20 codes (i.e., $2\times 10$ through 2-dimensional Softmax coding at each level), enabling a reduction of up to $O(\log N)$. Building upon this concept, Hier-SLAM [25] was first introduced as an RGB-D Gaussian Splatting SLAM that employs a hierarchical tree to represent semantic information in an unknown environment, achieving strong performance with improved storage efficiency and operational speed.
@@
Hier-SLAM [25] have demonstrated the advantages and potential of incorporating hierarchical categorization for a semantic RGB-D SLAM system. However, it is limited by a simplistic hierarchical tree representation and lacks robust semantic optimization. To address these limitations and further improve its effectiveness, we propose Hier-SLAM++, a comprehensive neuro-symbolic semantic Gaussian Splatting SLAM framework that incorporates an advanced hierarchical representation for semantic information. Different from Hier-SLAM, we incorporate both geometric (i.e., shape and size details) and semantic information into the hierarchical symbolic tree expression using both Large Language Models (LLMs) and 3D Generative models. Specifically, we utilize a text-to-3D generative model [26] to extract shape information for each class and organize it into hierarchical levels. Additionally, we leverage LLMs to include both semantic and size information in the tree structure, effectively addressing the lack of size information in 3D generative models. With the established tree, semantic information in the environment is represented as symbolic nodes along root-to-leaf paths, to be learned in an end-to-end manner during the SLAM operation. To enhance semantic understanding, we introduce an improved hierarchical loss function that incorporates both inter-level and cross-level optimizations. This coarse-to-fine hierarchical approach based on the established hierarchical representation is beneficial for real-world applications, especially those involving observations from distant to nearby views. Furthermore, our SLAM system is extended to support both RGB-D and monocular inputs. We utilize the geometric output of the 3D feed-forward model [27] as a geometric prior, removing the depth dependency of existing SLAM methods. This enables Hier-SLAM++ to perform semantic SLAM without relying on dedicated depth sensors, broadening the range of applicable scenarios. Finally, we enhance and refine the 3D Gaussian Splatting SLAM to improve both performance and efficiency. We conduct experiments on multiple synthetic and real-world datasets. The results demonstrate that our SLAM system outperforms or is on par with state-of-the-art NeRF-based and Gaussian-based SLAM methods. Additionally, our approach effectively reduces storage and training time.
The technical contributions of our work include:
@@
1) We introduce an advanced hierarchical tree representation that integrates semantic, shape, and size information, leveraging the capabilities of both LLMs and 3D generative models. This tree-based coding efficiently compresses semantic information while maintaining its inherent hierarchical structure, resulting in reduced memory usage and shorter training times.
2) We present a novel optimization approach for the proposed semantic hierarchical representation, integrating both inter-level and cross-level losses to ensure thorough optimization across all levels of the hierarchical semantic information.
@@
3) We propose an improved semantic SLAM system capable of supporting both RGB-D and monocular inputs by leveraging a 3D feed-forward model. To the best of our knowledge, Hier-SLAM++ is the first Gaussian Splatting based semantic SLAM system to support monocular input, significantly reducing sensor requirements and broadening its application scope.
4) We conduct extensive experiments on both synthetic and real-world datasets, comprehensively demonstrating the effectiveness of our method and its improvements.
@@ II-A Visual SLAM system
Visual Simultaneous Localization and Mapping (SLAM) using visual sensors is a long-standing and active research topic that focuses on the simultaneous estimation of ego-motion and the reconstruction of a global 3D map. Traditional SLAM methods [28, 29, 30, 31], developed over many years, have demonstrated both high accuracy and robustness, even in challenging environments. This makes them an indispensable frontend foundation for various real-world applications, such as autonomous driving and drones. In recent years, several new 3D representations have emerged in the 3D computer vision fields. With the advent of neural implicit representations, numerous works [32, 33, 34, 35] have demonstrated the effectiveness of NeRF-based SLAM approaches. These new representations introduce novel capabilities to the SLAM community, including enhanced rendering capability. 3D Gaussian Splatting is the latest advancement in the 3D vision field, offering significantly faster rendering speed and strong performance compared with neural implicit representations. Recent works [16, 17, 20, 19] have demonstrated new SLAM capabilities by integrating 3D Gaussian Splatting techniques, further pushing the boundaries of real-time scene understanding and reconstruction.
@@ II-B Neural implicit visual SLAM
Neural Radiance Fields (NeRF) [36] is a neural rendering technique that represents 3D scenes using a fully connected neural network. iMap [32] proposes a real-time keyframe-based RGB-D SLAM method that uses a multilayer perceptron (MLP) as the sole scene representation. Nice-SLAM [33] introduces a dense SLAM system that incorporates multi-level local information by employing a hierarchical scene representation. Nicer-SLAM [34] extends this approach by proposing a monocular SLAM system with a hierarchical neural implicit map representation, eliminating the dependency on depth information. Integrating SLAM with semantic understanding, many works [7, 11, 12] have explored neural implicit representations for semantic mapping and localization tasks. DNS-SLAM [7] incorporates 2D semantic priors with a coarse-to-fine geometry representation to embed semantic information into the map. SNI-SLAM [11] integrates geometry, appearance, and semantic features into a shared feature space to enhance the robustness of the SLAM method. However, these approaches are constrained by the slow convergence of neural implicit map representations, which often leads to performance degradation when coupled with semantic objectives [37, 38]. In contrast, Gaussian Splatting provides significant advantages, including fast rendering performance and high-density reconstruction quality, making it a promising alternative for map representation.
@@
3D Gaussian Splatting has recently gained attention as a promising 3D representation due to its fast rendering performance and high-density reconstruction quality. Leveraging this 3D representation, SplaTAM [16] uses silhouette guidance for pose estimation and map reconstruction in RGB-D SLAM systems, while MonoGS [17] supports both monocular and RGB-D SLAM. Photo-SLAM [19] utilizes 3D points generated by ORB-SLAM as position parameters for Gaussian Splatting primitives within the SLAM system. These studies highlight the strong potential of 3D Gaussian Splatting in SLAM tasks across various scenarios [16, 17, 18, 19, 20]. However, integrating semantic understanding into SLAM presents a significant challenge, as it requires jointly solving three high-dimensional optimization problems with varying value ranges and convergence behaviors, while also significantly increasing storage requirements. Focus on the integration of semantic information into the Gaussian Splatting SLAM, SGS-SLAM [23] added RGB 3-channels to create a semantic visualization map to avoid true semantic labels learning. SemGauss-SLAM [24] uses a flat semantic representation supervised by a large pre-trained foundation model. However, these methods overlook the natural hierarchical structure of the real world. Additionally, relying on large foundation models increases the complexity of the neural network and its computational demands, with performance largely dependent on the embeddings from these pre-trained models. Eliminating the dependency on foundation models, this work proposes a simple yet effective hierarchical representation for semantic understanding in the SLAM system which largely reduces the training storage and time.
@@ III Method
Figure 2: Left: Overview of the Hier-SLAM++ pipeline. The global 3D Gaussian map is initialized using the first frame of the video stream input. The system then alternates between the Tracking and Mapping steps as new frames are processed (see Section III-D). In the RGB-D setting, depth is directly obtained from sensor input, whereas in the monocular setting the 3D feed-forward method (DUSt3R) is used to generate a geometric prior for depth estimation (see Section III-C). Top Right: Hierarchical representation of semantic information. The Tree Generation process leverages the capabilities of both LLMs and 3D Generative Models in a top-to-bottom manner. This hierarchical tree is used to establish a symbolic coding for each Gaussian primitive (see Section III-A). Additionally, we introduce a novel loss function that combines Inter-level Loss $L_{\text{Inter}}$ and Cross-level Loss $L_{\text{Cross}}$ to optimize the hierarchical semantic representation (see Section III-B). Bottom Right: An example of hierarchical semantic rendering.
The entire pipeline of our method is illustrated in Fig. 2. Within Hier-SLAM++, we model the semantic space as a hierarchical tree, where semantic information is represented as a root-to-leaf hierarchical symbolic structure. This representation not only comprehensively captures the hierarchical attributes of semantic information, but also effectively compresses storage requirements, while enhancing semantic understanding with valuable scaling-up capability. To construct an effective tree, we leverage both LLMs and 3D Generative Models to integrate semantic and geometric information. Next, each 3D Gaussian primitive is augmented with a hierarchical semantic embedding code, which can be learned in an end-to-end manner through the proposed Inter-level and Cross-level semantic losses. Furthermore, our method is extended to support monocular input by incorporating geometric priors extracted from the feed-forward method, thereby eliminating the dependency on depth information. The following sections provide detailed descriptions for each part of our Hier-SLAM++.
III-A Hierarchical representation
We propose hierarchical symbolic representation for semantic information to provide improved efficiency, compression, and compactness for semantic SLAM, making it well-suited for multi-parameter joint optimization and enabling scalability to complex scenes. In this section, we provide a detailed explanation of the hierarchical tree parametrization and its generation process.
III-A1 Tree Parametrization
The hierarchical tree $\bm{G}$ adopted in Hier-SLAM++ is represented as the set of vertices and edges, $\bm{G}=(\bm{V},\bm{E})$. The vertices set $\bm{V}=\cup_{l=0}^{L}\left\{\bm{v}_{l}\right\}$ comes from the classes from all tree levels, where $\left\{\bm{v}_{l}\right\}$ represents the set of nodes at the $l$-th level of the tree, and we use $L$ to represent depth of the whole tree. The edge set $\bm{E}=\cup_{m=0}^{L-1}\left\{\bm{e}_{m}\right\}$ captures the subordinate relationships, reflecting semantic attribution, size message, and geometric knowledge. For the $i$-th semantic class $g^{i}$, which is treated as a single leaf node in the tree view, its hierarchical expression is:

### v2

@@ Abstract
We propose Hier-SLAM++, a comprehensive Neuro-Symbolic semantic 3D Gaussian Splatting SLAM method with both RGB-D and monocular input featuring an advanced hierarchical categorical representation, which enables accurate pose estimation as well as global 3D semantic mapping. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making scene understanding particularly challenging and costly. To address this problem, we introduce a novel hierarchical representation that encodes both semantic and geometric information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs) as well as the 3D generative model. By utilizing the proposed hierarchical tree structure, semantic information is symbolically represented and learned in an end-to-end manner. We further introduce an advanced semantic loss designed to optimize hierarchical semantic information through both Intra-level and Inter-level optimizations. Additionally, we propose an improved SLAM system to support both RGB-D and monocular inputs using a feed-forward model. To the best of our knowledge, this is the first semantic monocular Gaussian Splatting SLAM system, significantly reducing sensor requirements for 3D semantic understanding and broadening the applicability of semantic Gaussian SLAM system. We conduct experiments on both synthetic and real-world datasets, demonstrating superior or on-par performance with state-of-the-art methods, while significantly reducing storage and training time requirements. Our project page is available at: https://hierslampp.github.io/.
@@ I Introduction
Visual Simultaneous Localization and Mapping (SLAM) is a key technology for ego-motion estimation and scene perception, and is widely employed across various domains such as autonomous drones [1], self-driving vehicles [2], and interactive experiences in Augmented Reality (AR) and Virtual Reality (VR) [3]. Semantic information, which provides high-level knowledge about the environment, is indispensable for comprehensive scene understanding. It plays a crucial role in enabling intelligent robots to perform complex tasks. In recent decades, advances in image segmentation and enriched map representations have significantly accelerated progress in semantic visual SLAM [4, 5, 6, 7, 8, 9, 10, 11, 12].
@@
Recently, 3D Gaussian Splatting has emerged as a powerful representation for 3D scenes [13, 14, 15], owing to its fast rendering and optimization efficiency enabled by the highly parallelized rasterization of 3D primitives. Specifically, it models the continuous distribution of geometric attributes using Gaussian functions. This formulation not only enhances the 3D representation fidelity but also facilitates efficient optimization, making it particularly suitable for SLAM tasks. While promising, current SLAM systems based on 3D Gaussian Splatting [16, 17, 18, 19, 20] primarily focus on geometric reconstruction. However, the absence of semantic integration limits their applicability in complex downstream tasks that require high-level scene understanding.
To enable semantic understanding in SLAM systems, a straightforward approach, analogous to the geometric modeling of 3D Gaussians, is to augment each 3D primitive with a discrete semantic label and represent its distribution using a categorical distribution, typically implemented via a flat Softmax-based representation. However, 3D Gaussian Splatting is already a storage-intensive approach [21, 22], requiring numerous primitives with rich parameters to achieve high-quality rendering. Adding semantic distributions for each primitive further increases memory and computational costs, with complexity growing linearly with the number of semantic classes, making it impractical for complex semantic understanding. Moreover, given a limited number of observations (e.g., video frames), introducing high-dimensional semantic parameters leads to the curse of dimensionality, hindering optimization performance. This added overhead is especially problematic in robotics, where computational efficiency and resource constraints are critical. Therefore, a key challenge is to design semantic encoding strategies that are both efficient and scalable, enabling robust semantic SLAM without compromising runtime and memory efficiency.
Semantic information often exhibits a natural hierarchical structure. As illustrated in Fig. 1, semantic classes can be abstracted into multi-level concepts, such as structures and surfaces, which can be further organized into a hierarchical symbolic tree. With careful design, this structure enables rich semantics to be encoded efficiently using a small number of symbolic nodes, yielding a compact and scalable representation. For instance, a binary tree of depth $10$ can represent up to $2^{10}$ classes using only $20$ nodes (i.e., $2\times 10$ nodes through 2-dimensional Softmax at each level), reducing the complexity from $O(N)$ to $O(\log N)$.
@@
Building upon this concept, we propose Hier-SLAM++, a comprehensive neural-symbolic semantic SLAM framework based on 3D Gaussian Splatting, featuring a novel hierarchical representation for semantic information. To design an effective hierarchical structure, we incorporate both semantic and geometric cues, including shape and size information. Specifically, we leverage a text-to-3D generative model [23] to extract shape-aware features for each class and organize them into hierarchical levels. Additionally, we employ Large Language Models (LLMs) to enrich the hierarchy with semantic relations and size priors, compensating for the limitations of generative models. With the constructed tree, semantic information in the environment is represented as symbolic nodes along a hierarchical path and learned end-to-end during SLAM optimization. Specifically, we propose two types of hierarchical semantic representations, one-hot and binary semantic embedding, which offer different degrees of compactness. To enhance semantic understanding, we introduce a novel hierarchical loss function that performs joint optimization at multiple levels, including Intra-level and Inter-level supervision. Furthermore, Hier-SLAM++ supports both RGB-D and monocular settings. We employ the geometric output of a 3D feed-forward model [24] as a depth prior, effectively removing the reliance on dedicated depth sensors. This enables monocular semantic SLAM and broadens the applicability of our system. We further refine the 3D Gaussian Splatting-based SLAM pipeline to improve both tracking performance and runtime efficiency. Experiments on both synthetic and real-world datasets demonstrate that Hier-SLAM++ achieves superior or on-par performance with state-of-the-art methods in localization, mapping, and semantic understanding under the RGB-D setting, while maintaining competitive performance in the monocular setting. The main contributions are as follows:
@@
1) We propose a hierarchical tree structure that integrates semantic, shape, and size information, constructed using LLMs and 3D generative models. This representation efficiently encodes semantic messages while preserving its structural hierarchy, leading to reduced memory consumption and faster training.
2) Based on the generated hierarchical tree, we propose two types of semantic embeddings that offer different degrees of compactness. Furthermore, we introduce a novel optimization strategy for hierarchical semantic representations, incorporating both Intra-level and Inter-level losses to enable comprehensive refinement of both embeddings.
@@
3) We develop an enhanced SLAM framework that supports both RGB-D and monocular settings by leveraging a 3D feed-forward model as a geometric prior. To the best of our knowledge, Hier-SLAM++ is the first Gaussian Splatting-based semantic SLAM system capable of operating with monocular input, significantly lowering sensor requirements and improving deployment flexibility.
This paper builds upon our previous work [25], with the following major extensions: 1) The tree construction process is enhanced by integrating 3D generative models and improving the use of LLMs, resulting in a more effective symbolic tree structure. 2) In addition to the one-hot semantic representation from our earlier work, we propose a binary semantic representation for further compactness. 3) We improve semantic optimization with a new hierarchical loss formulation and introduce a comprehensive hierarchical optimization strategy applicable to both one-hot and binary representations. 4) We propose a monocular pipeline to enable robust monocular SLAM by leveraging 3D feed-forward models, thereby extending the RGB-D semantic SLAM system into a unified framework that supports both RGB-D and monocular inputs. Additionally, we include new real-world datasets and experiments to further validate the proposed approach.
@@ II-A Visual SLAM system
Visual Simultaneous Localization and Mapping (SLAM) is a long-standing topic in computer vision and robotics. It jointly estimates camera poses and reconstructs global maps by optimizing appearance-based objectives [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. One classical monocular-based approach aims to estimate accurate camera poses and a sparse global map, by minimizing the reprojection error terms [28, 29, 30], or by minimizing the photometric losses [31, 32, 33]. DTAM [34] takes a different direction by performing dense reconstruction and camera tracking with RGB inputs, though real-time capability relys on GPU support. Another category of methods adopts RGB-D input for dense 3D reconstruction, which is able to provide richer information. The works [35, 36] jointly estimate camera poses and dense surfaces from RGB-D input using a volumetric TSDF map. Alongside advances in purely geometric SLAM, semantic SLAM has gained increasing attention for its ability to provide high-level scene understanding [37, 38, 39, 40, 41, 42, 43, 44, 8, 6]. Sparse semantic SLAM methods rely on pre-scanned models [37, 38] or use bounding boxes [39, 40, 41] and quadrics [42], though accuracy remains limited. In contrast, semantic understanding in dense SLAM becomes more tractable and effective when combined with dense observations from RGB-D inputs [45, 46, 47] or dense visual SLAM pipelines [44]. More recently, new paradigms have emerged to enhance dense SLAM with richer scene understanding across both monocular and RGB-D inputs, such as implicit neural fields and 3D Gaussian Splatting. Furthermore, semantic integration allows current SLAM systems to move beyond geometric reconstruction to deliver high-level scene interpretation, such as class-level mapping, semantic navigation, and task-aware perception. This significantly broadens its applications in real-world fields.
Some works, such as Kimera [48], adopt dynamic scene graphs for representing entities within the observed environment  [49, 50, 51, 52, 53]. In contrast, we propose a generalizable hierarchical symbolic representation for modeling semantics in the world, leveraging large language models and 3D generative priors, along with symbolic capabilities [54, 55] to abstract semantic concepts without relying on scene-specific definitions. Crucially, our symbolic hierarchy is carefully designed to produce compact semantic embeddings and is seamlessly integrated into the neural learning process within the SLAM system, offering both structural abstraction and interpretability for semantic encoding.
@@ II-B Neural implicit visual SLAM
Neural Radiance Fields (NeRF) [56] introduce a novel scene representation paradigm, inspiring the integration of NeRF with SLAM pipelines for simultaneous localization and high-quality mapping. Unlike traditional SLAM methods that rely on explicit geometry, NeRF-SLAM models represent scenes implicitly, enabling continuous and photorealistic reconstruction from sequential views. Early work such as iNeRF [57] explored solely pose tracking with a fixed pre-trained neural field. iMap [58] advanced this idea by proposing a keyframe-based RGB-D SLAM that jointly optimizes camera poses and a global implicit map using a single MLP. NICE-SLAM [59] enhances robustness and scalability by introducing a hierarchical neural implicit representation, which is later extended to the monocular setting in NICER-SLAM [60]. Recent works focus on improving efficiency: ESLAM [61] leverages a hybrid representation combining implicit TSDFs with multi-scale feature planes to achieve faster operation; Point-SLAM [62] adopts a point-based neural representation to enable efficient and detail-preserving dense SLAM. To further enhance scene understanding beyond geometry and appearance, researchers also incorporate semantics into NeRF-based SLAM systems [7, 11, 12, 63]. Some works leverage the semantic information extracted from 2D observations to integrate them into the visual SLAM system and generate the global semantic map [7, 63]. SNI-SLAM [11] proposes a semantic RGB-D SLAM which learn semantics by integrating geometry, appearance, and semantic features into a shared feature space. However, these methods remain constrained by the slow convergence and high overhead of neural implicit representations [64, 65], especially when jointly optimizing semantics. The inherent inefficiency of NeRF representations still conflicts with the real-time demands of SLAM systems—let alone the added cost from semantic integration. In contrast, Gaussian Splatting provides significant advantages, including fast rendering performance and high-density reconstruction quality, making it a promising new kind of map representation.
@@
3D Gaussian Splatting has recently emerged as a promising 3D representation due to its fast rendering and high-fidelity reconstruction. Compared to implicit representations, it models the environment as a collection of explicit Gaussian distributions with geometric and appearance attributes. Several SLAM systems have adopted this representation in their pipelines. SplaTAM [16] proposes a RGB-D SLAM framework with isotropic Gaussians representation using silhouette guidance for pose estimation, while MonoGS [17] presents an efficient solution supporting both monocular and RGB-D inputs. Photo-SLAM [19] uses 3D points from ORB-SLAM to initialize Gaussian positions; GS-SLAM [18] introduces a coarse-to-fine tracking strategy with sparse Gaussian selection; and Gaussian-SLAM [20] utilizes DROID-SLAM poses to manage multiple 3D Gaussian submaps. These approaches showcase the versatility of 3D Gaussian Splatting in various SLAM settings. However, integrating semantics into Gaussian Splatting-based SLAM remains challenging. Jointly optimizing geometry, appearance, and semantics involves heterogeneous value ranges and significantly increases memory and computational demands. GS3LAM [66] fuses 2D semantic labels into an RGB-D SLAM framework, with embeddings sized according to the number of semantic classes. SGS-SLAM [67] appends RGB 3-channels to support semantic visualization. SemGauss-SLAM [68] leverages flat semantic embeddings supervised by foundation models, though this introduces high computational cost and overlooks the hierarchical structure of real-world semantics. To address these limitations, we propose a compact and efficient solution that fully exploits the hierarchical nature of semantic categories. Our method encodes semantic information into a hierarchical symbolic tree, generated through the synergy of large language models and 3D generative priors, and learns this semantic representation end-to-end within the SLAM framework operation.
@@ III Method
Figure 2: Overview of the Hier-SLAM++ pipeline. Top Left: Hierarchical representation of semantic information. The Tree Generation process leverages the capabilities of both LLMs and 3D Generative Models (see Section III-A). This hierarchical tree is used to establish a symbolic coding for each Gaussian primitive. Additionally, we introduce a novel loss function that combines Intra-level Loss $L_{\text{Intra}}$ and Inter-level Loss $L_{\text{Inter}}$ to optimize the hierarchical semantic representation (see Section III-B). Bottom Left: An example of hierarchical semantic rendering. Right: The global 3D Gaussian map is initialized using the first frame of the video stream input. The system then alternates between the Tracking and Mapping steps as new frames are processed (see Section III-D). In the RGB-D setting, depth is directly obtained from sensor input, whereas in the monocular setting the 3D feed-forward method (DUSt3R) is used to generate a geometric prior for depth estimation (see Section III-C).
The overall pipeline of our method is illustrated in Fig. 2. At the core of our approach, we model the entire semantic space as a compact hierarchical structure, where each semantic class is represented by a symbolic hierarchy. To construct a generative hierarchical tree that captures both semantic relations and associated geometric attributes, we leverage Large Language Models (LLMs) and 3D generative models. The hierarchical representation and tree generation process are detailed in Section III-A. Based on the proposed structure, semantic information is encoded into the embedding space in two formats: a one-hot representation and a binary representation, both of which are hierarchically organized and compact. To support end-to-end training, we introduce hierarchical optimization with two loss functions, Intra-level and Inter-level, which jointly optimize the semantic space based on incoming observations. The two proposed tree encoding schemes and hierarchical optimization are discussed in detail in Section III-B. In the monocular setting, we leverage geometric priors extracted from a feed-forward 3D model to guide learning during SLAM, incorporating online depth adjustment and geometric supervisory signals, as described in Section III-C. Finally, the full semantic Gaussian Splatting SLAM system is presented in Section III-D.
III-A Hierarchical Representation and Generation
Leveraging the inherent hierarchy of semantic classes, we propose a hierarchical symbolic representation to encode semantic information in a compact and efficient manner, enhancing the scalability of semantic SLAM system. This section presents a detailed description of the proposed hierarchical tree parametrization and its generation process using both LLMs and 3D generative models.
III-A1 Hierarchical Tree Parametrization
The proposed hierarchical representation is abstracted as a multi-level tree, formulated as $\bm{G}=(\bm{V},\bm{E})$, where $\bm{V}$ and $\bm{E}$ denote the sets of vertices and edges, respectively, as illustrated in Fig. 3. The vertex set is defined as $\bm{V}=\left\{v_{l}\right\}_{l=0}^{L}$, where $\bm{v}_{l}$ represents the set of semantic classes at level $l$ of the tree hierarchy, spanning from the root level ($l=0$) to the deepest level ($l=L$). The edge set $\bm{E}=\left\{e_{m}\right\}_{m=0}^{L-1}$ encodes the subordinate relationships between adjacent levels in the hierarchy, capturing semantic attribution, relative scale, and geometric structure. The original semantic classes are represented as leaf nodes located at the deepest level ($l=L$) of the hierarchical tree. For the $i$-th semantic class $g^{i}$, its hierarchical representation is defined as: